An environmental trust will pay $73,726 once every 3 years for remediation of a site for an indefinitely long period of time. The interest rate is 7% per year. If the project is due to start, and the first payment made, in 3 years, how much money must be invested today? Round to the nearest dollar.
Answer #1

ANSWER:

I = 7% AND N = 3 YEARS

PV = Amount paid every 3rd year(p/f,i,n) + Amount paid in 6th year(p/f,i,n) + ...... + Amount paid in infinitely(p/f,i,n)

pv = 73,726(p/f,7%,3) + 73,726(p/f,7%,6) + .... + 73,726(p/f,7%,infinity)

pv = 73,726(p/f,7%,3) * 6.6686 (6.6686 is the total of all the p/f factors till infinity)

pv = 73,726 * 0.8163 * 6.6686

pv = 327,608.09 or $327,608

so the present value is $327,608